Tucker Corporation plans to issue new 20-year bonds that are callable after 5 years at a call premium of $1,050. Suppose this bond has a face value of $1,000 the price is currently $1,000. The coupon rate is 10% which is paid semiannually. Does the yield to maturity exceed the yield to call on this bond?

a. Yes

b. No

c. Not enough information
Answer #1

Yield to maturity:

The yield to maturity will be same as coupon rate since price is equal to face value

Yield to maturity = 10%

Yield to call:

Number of periods = 5 * 2 = 10

Coupon = (0.1 * 1000) / 2 = 50

Yield to call = 10.78%

Keys to use in a financial calculator: 2nd I/Y 2, FV 1050, PV -1000, PMT 50, N 10, CPT I/Y

NO yield to maturity does not exceed the yield to call on this bond